### v1

# Building Age Estimation: A New Multi-Modal Benchmark Dataset and Community Challenge

Abstract
Estimating the construction year of buildings is of great importance for sustainability. Sustainable buildings minimize energy consumption and are a key part of responsible and sustainable urban planning and development to effectively combat climate change. By using Artificial Intelligence (AI) and recently proposed Transformer models, we are able to estimate the construction epoch of buildings from a multi-modal dataset. In this paper, we introduce a new benchmark multi-modal dataset, i.e. the Map your City Dataset (MyCD), containing top-view Very High Resolution (VHR) images, Earth Observation (EO) multi-spectral data from the Copernicus Sentinel-2 satellite constellation, and street-view images in many different cities in Europe, co-localized with respect to the building under study and labelled with the construction epoch. We assess EO generalization performance on new/ previously unseen cities that have been held-out from training and appear only during inference. In this work, we present the community-based data challenge we organized based on MyCD. The AI4EO Challenge MapYourCity was opened in 2024 for 4 months. We present the Top-4 performing models, and the evaluation results. During inference, the performance of the models using both all three input modalities and only the two top-view modalities, i.e. without the street-view images, is examined. The evaluation results show that the models are effective and can achieve good performance on this difficult real-world task of estimating the age of buildings, even on previously unseen cities, as well as even using only the two top-view modalities (i.e. VHR and Sentinel-2) during inference.
I Introduction
Estimation of the age of buildings in cities is important for sustainability, urban planning, and structural safety purposes. Sustainable buildings minimize energy consumption and are a key part of responsible and sustainable urban planning and development that seeks to effectively combat climate change, as well as rising temperatures. With increase in urbanization and Earth Sciences research into urban areas, the need to estimate the energy efficiency of buildings, so as to meet important sustainability goals, has never been greater. The age of buildings plays an important role in energy modelling, as well as in urban policies. The age of the building can be used as a proxy for the energy efficiency. However, data for the construction period of buildings is not currently available globally in a consistent and homogeneous manner. To this end, as a start, in this study, we focus on European cities.
In this real-world application study for sustainability and climate, the main aim is to develop models that automatically estimate the construction year of buildings with improved performance using Artificial Intelligence (AI) techniques and cross-view datasets based on street-view images and Earth Observation (EO) datasets. The goals of this paper are to present the new multi-modal dataset, examine example images from this street-view and EO dataset, present the community data challenge that we organized, and present the evaluation results of the top-performing models from the challenge. We introduce a new benchmark dataset for a real-world multi-modal application study. It consists of three modalities, including street-view ground images, top-view satellite Very High Resolution (VHR) images at 50 cm resolution, and Copernicus Sentinel-2 multi-spectral satellite data. All the modalities are co-localized with respect to the specific building under study, as well as its label of the construction epoch. The problem is formulated as a classification task with seven classes for the construction epoch, ranging from 1900 to nowadays.
This paper evaluates the EO generalization performance of the models on cities that are not included in the training dataset, i.e. on new/ previously unseen cities. We examine the performance of the models when both training and testing are performed with all 3 modalities. There is benefit in using the top-view VHR images (and to a lesser extent, Sentinel-2 data), in addition to the street-view images. In this work, we also examine the performance of the models when training is performed with all the three modalities, and testing is performed without the street-view images, i.e. when using only two out of the three encoders (i.e. top-view VHR and Sentinel-2 images). These results are compared to those when training and testing are performed using all the three modalities.
Figure 1: Left: Example street-view image from the multi-modal dataset. Right: Example top-view satellite VHR image.
Main contributions. Our contributions are that this work introduces a novel multi-modal dataset, presents the challenge we organized to estimate the age of buildings in different cities in Europe using this dataset, examines the performance evaluation results of the models from the challenge, and shows the potential of effectively combining and fusing street-view, top-view satellite and remotely sensed Sentinel-2 EO data for accurately predicting the construction period of buildings. In this paper, we present and compare the performance of the models from the challenge that have been trained and tested on the new multi-modal street-view and EO dataset.
II Data and Task
II-A The multi-modal street-view and EO dataset MyCD
We present the multi-modal dataset and examine example images from this street-view and EO dataset. The dataset MyCD encompasses the three modalities of: a) Street-view images, b) Top-view satellite VHR RGB images at 50 cm resolution, and c) Multi-spectral 12-band S-2 data. These input modalities have very different characteristics (number of spectral bands, spatial resolution, and building perspectives from top and street-view). In the dataset, the three modalities are co-localized with respect to the specific building under study. In Fig. 1, we show two example images for the modalities of street-view ground and top-view satellite VHR, respectively.
Figure 2: Example top-view satellite VHR image and the corresponding geo-localized S-2 image (here, RGB bands for visualization purposes, where the dataset includes all the $12$ spectral bands, i.e. L2A) from the multi-modal dataset MyCD.
In Fig. 2, an example top-view VHR image and the corresponding S-2 image, both from the new multi-modal dataset MyCD, are shown, together with the building location in the centre, as a blue dot in the image. We observe the geospatial overlay of the two top-view image modalities and that the two top-view modalities are centered at the building under study.
In Fig. 3, additional examples of street-view and top-view images from the new multi-modal dataset MyCD are shown.
The dataset MyCD comprises cross-view image data from $6$ countries in Europe, and more specifically from $19$ cities, for $7$ different building construction epoch classes extracted from the EUBUCCO database. For the training set data, $15$ cities have been chosen, while for the test set data, the remaining $4$ cities have been used. The building construction epoch classes are $7$, i.e. labelled from 00 to $6$, where Class 00 is the building construction age “Before 1930”, Class $1$ the period 1930-1945, Class $2$ the epoch 1946-1960, Class $3$ the period 1961-1976, Class $4$ the time epoch 1977-1992, Class $5$ the period 1993-2006, and Class $6$ “After 2006”. In Figs. 4 and 5, we show the number of image samples in the dataset MyCD per country and building construction epoch class, as well as per city.
II-B The task of predicting the age of buildings
The multi-modal dataset MyCD was used in the challenge that we organized in 2024 to predict the age of buildings in several different cities in Europe. The goal of the challenge was to achieve the best performance possible on cities not included in the training data, i.e. on previously unseen cities. In addition, half of the test samples did not include the street-view modality. With this data challenge, we hope to advance the research on models’ out-of-distribution (OoD) generalization.
We present the evaluation results for the OoD generalization performance of the top-performing models from the challenge when: 1) both training and testing are performed with all the three modalities of the dataset MyCD, and 2) training is performed with the three modalities of the street-view and EO dataset and testing/ inference is performed using only the top-view satellite modalities, i.e. without the street-view images.
Figure 3: Left: Example street-view image of building from the dataset MyCD, and the corresponding top-view VHR image. The images are co-localized with respect to the specific building. Right: Example top-view image from our multi-modal dataset.
Figure 4: Number of images per country and building age class.
Figure 5: Number of images per city and building age class.
III Models and Methodologies
III-A Models developed to predict building construction year
In this section, we will discuss the models from the challenge that estimate the age of buildings. More specifically, we discuss the models that achieve the top/ winning results of the challenge (i.e. in Sec. V, in Figs. 11 and 12). In particular, in this section, we will discuss four different models that predict the construction epoch of buildings. To achieve very good/ top performance, they employ several different deep learning techniques, for example class label correlation modelling.
Label correlation. The team Creamble performs label correlation modelling to take advantage of the distance between the labels. For the seven building construction epoch classes, adjacent classes are correlated, while far apart classes are not. Soft labels are also used, as well as label smoothing and hyperparameter search. In addition, K-fold cross-validation, the recently proposed model EVA [1], and a model ensemble of $15$ models with soft-voting are also used. The model minimizes the cross-entropy loss with label smoothing (0.3). The flowchart is shown in Fig. 6, where the model ensemble that estimates the construction period of buildings is illustrated.
Figure 6: Flowchart of the model to predict the age of buildings by the team Creamble.
Figure 7: Flowchart diagram of the model by the team TelePIX AI, where the results are in Sec. V, in Figs. 11 and 12.
Prediction-level data fusion. The team TelePIX AI performs prediction-level data fusion (rather than data fusion at the feature level), as well as K-fold cross-validation with the city as the fold for each country. For this method, a mean confidence model ensemble is used. Moreover, as encoders, the model EVA-02 [2] is also utilized, as well as reBEN [3]. For classification, the cross-entropy loss is used. The flowchart diagram of this model is illustrated in Fig. 7. Hyper-parameter optimization using Bayesian optimization with Optuna (e.g. for the learning rate (lr), batch size, and lr decay factor) is also performed. In addition, here, the image pre-processing is: normalization, histogram equalization, random rotation and flipping, and random hide of the street-view images, where the latter is for the missing modality problem during inference (i.e. missing street-view images). For this approach, the key factors for good model performance are: powerful encoders, random hide street-view (P: 0.5) so that the model learns the significant features from the VHR images and Sentinel-2, split based on country, ensemble technique and pre-processing to increase robustness of the final prediction and training stability.
Figure 8: Flowchart of model using an additional shared feature encoder for the street- and top-view VHR images and the distribution alignment loss, by the team xmb.
Figure 9: Flowchart of the model that uses SwinT encoders, by the team The AI Buzzard.
Ensemble with country categorical feature. The team xmb trains an ensemble of four models. For this approach, a categorical feature for the country in the format of one-hot representation is used. This country categorical feature is used in the late data fusion concatenation of the features. For the three input modalities of the dataset MyCD, the latent representations after average pooling are combined, also using the country categorical feature, and then inputted to the final dense Fully Connected (FC) layer. In addition, in the ensemble, the models EfficientNet and MobileViT are also used, and geometric data fusion is also performed. The latent feature representations before average pooling are combined according to the method presented in [4], and then inputted to average pooling and, then, concatenated using the country categorical feature. To address the problem of the missing modality during inference, street-view images during training are randomly replaced with all zeros, i.e. input dropout, and shared feature modelling [5] is performed. Here, an additional shared encoder for the two different modalities of street-view and top-view VHR is used. The flowchart diagram is shown in Fig. 8. The distribution alignment loss function term [5] is also added in the objective function minimized during training.
SwinT-encoder late data fusion with attention. The model Swin Transformer (SwinT) pretrained on ImageNet22K is used by the team The AI Buzzard, for the encoders of the three different input modalities. Late data fusion using attention is performed, as well as 5-fold cross-validation training, grouped by cities and stratified by the class label. Data fusion is performed by computing a weighted sum of the embeddings. In this way, the model is able to handle missing data during inference (i.e. missing street-view images). The flowchart diagram is shown in Fig. 9. For this method, the models are first trained on all the data and, then, they are fine-tuned on the individual countries. For data augmentation, random flipping, color jittering, and random resized cropping are utilized.
Discussion about the key features. The main characteristics of the models discussed in the preceding paragraphs are: i) very good/ top pretrained model, i.e. EVA [1] or EVA-02 [2], that is used as a starting point so as to not train from scratch, ii) model ensemble, iii) hyper-parameter optimization using for example Optuna, iv) label correlation modelling between adjacent classes (per modality) because the distance between adjacent classes is smaller than the distance between classes that are not adjacent, and v) missing modality distribution alignment loss together with a shared feature encoder. Here, additional key features of the main models are: K-fold cross-validation training, geometric data fusion using the methodology in [4], and use of a categorical feature for the country in the form of one-hot vector representation.
III-B Baseline model: Age of Buildings for Sustainability (ABS)
We present the baseline model Age of Buildings for Sustainability (ABS) which is based on Transformers, and more specifically on the recently proposed model SegFormer [6, 7], which uses multi-scale features. This baseline model of the challenge employs three encoder networks, one for each of the modalities of the dataset MyCD, and performs late data fusion by concatenating the feature representations in the latent space. The flowchart of the model ABS is shown in Fig. 10. The output is the construction epoch class for the queried building. We also release our code for ABS for reproducibility [^1] The GitHub repository is: http://github.com/AI4EO/MapYourCity..
Figure 10: Flowchart of the model ABS, the baseline of the challenge we organized using the multi-modal dataset MyCD.
Figure 11: Evaluation results of the different models, and more specifically the private leaderboard of the challenge.
Figure 12: Results from the evaluation of the models, and more specifically the public leaderboard of the challenge.
IV Related Work
Many different AI models from the recent literature perform data fusion of street-view images and top-view satellite images. With AI and deep learning methods, by performing late data fusion (i.e. concatenating latent representations in the feature space), it is possible to achieve better performance than early data fusion, because street-view images and top-view satellite data have many differences, for example in resolution, object characteristics and scale, and significant differences in appearance. Street-view and top-view satellite images do not look alike and top-view data cover a larger geographical area. In general, the effective fusion of street-view ground images and nadir-view satellite or aerial images is nowadays an open research problem. By jointly using both street- and top-view images, the model proposed in [8] performs mapping of building functions using a specific data fusion strategy that combines in a decision-level manner different models trained for each image modality independently. In this way, by performing decision-level data fusion of the two modalities, the method presented in [8] combines the predictions of different models in order to accurately perform mapping of building functions, for example commercial, residential, public, and industrial, i.e. classification using four different classes.
Furthermore, nowadays, the importance of street-view images has become evident in the literature [9]. Street-view imagery is an important data source for urban analytics, helping us derive insights for making informed decisions [10]. Using street-view images, we can analyze for example the built environment, the vegetation, and the transportation [9], and the results of the analyses can be linked to important health, urban, and socio-economic studies. In this work, street-view images from Mapillary are used in the multi-modal dataset MyCD.
Street-view imagery is also important for urban infrastructure assessment [11] and urban mobility, as well as for spatial data infrastructures, urban planning [12], and urban greenery. In [11], the dataset Global Streetscapes is introduced which contains $10$ million street-view images from $688$ cities around the Earth, from $212$ countries and different regions. These images have been crowdsourced from Mapillary and KartaView. Moreover, the dataset Global Streetscapes [11] also includes metadata information for every street-view image, such as geo-location, longitude and latitude, acquisition time, and contextual, semantic, and perceptual information.
Transformer models have shown to achieve good/ top performance and outperform other models, for example models that use an architecture based on Convolutional Neural Networks (CNN), U-Net, and/or Residual Networks (ResNet). Importantly, in [13], geo-location matching of street-view images and top-view aerial images is performed using a Transformer-based model and the mechanism of attention. The model TransGeo proposed in [13] is based on Transformers because global information modeling is crucial. It performs attention-guided non-uniform cropping of images as a means of attending and zooming-in on only informative image patches, not focusing on non-informative (for the task) parts of the image.
Different datasets that contain top-view satellite or aerial images exist, for example the dataset OpenEarthMap [14]. This publicly available dataset is designed for the task of high-resolution land cover mapping semantic segmentation. More specifically, OpenEarthMap contains $5000$ aerial and satellite images, and for each image, it has land cover labels/ annotations for $8$ different classes. It covers in total $97$ regions, from $44$ countries across $6$ continents. This recent dataset OpenEarthMap has been used, for example, in the model proposed in [15] which performs semantic segmentation land cover classification in a few-shot continual learning setting.
V Evaluation and Results
Evaluation metrics. In this work, for the numerical evaluation results, we compute the classification confusion matrix for the task of predicting the construction period of buildings, where we have $7$ different classes, for example “After 2006”. We calculate the average of the diagonal items of the confusion matrix, and we report this metric, which is also known as the Mean Producer Accuracy (MPA) evaluation metric.
We present the evaluation of the models from the challenge in Figs. 11 and 12. We use MPA_ALL_MODALITIES to denote the setting when both training and testing are performed with all modalities. In addition, we also use MPA_TOP_VIEW to denote the setting when training is performed with all modalities and inference is performed using only the two top-view modalities. TOTAL is the average of the two MPAs.
In Figs. 11 and 12, in the results [^2] https://platform.ai4eo.eu/map-your-city/leaderboard, the MPA_TOP_VIEW scores are lower than MPA_ALL_MODALITIES. The top score for MPA_ALL_MODALITIES in the private leaderboard is $0.74$ indicating that the estimation of the construction epoch of buildings in unseen cities is possible. When using only the top-view data during inference, then the performance of the models decreases by more than $10$ points in the MPA score. Finally, we observe that for the TOTAL scores, the top winning values are approximately $0.65$ in the private leaderboard.
VI Conclusion
In this work, we have introduced the new multi-modal benchmark dataset MyCD and the real-world multi-modal application study for EO generalization. The performance evaluation of the models from the challenge show that estimating the age of buildings in previously unseen cities is possible and effective, as well as predicting the construction period of buildings/ houses from only top-view satellite images, i.e. without using the street-view images during inference.
Acknowledgment
The challenge was accomplished thanks to the contributions of many people. We want to thank all the team: Juan Pedro, Dennis Albrecht, and everyone in ESA AI4EO.